OpenStudy (anonymous):

Write the terms of the series and find their sum 4E (k+8)^2 K=1

OpenStudy (anonymous):

that E : \[\Sigma \]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\sum_{k=1}^4(k+8)^2\]

OpenStudy (anonymous):

YES that!

OpenStudy (anonymous):

replace \(k\) by 1 then by 2 then by 3, then by 4 put plus signs between them

OpenStudy (anonymous):

\[ \sum_{k=1}^4(k+8)^2 \]\(i=k+8\implies k=i-8\)\[ \sum_{i-8=1}^{i-8=4}(i)^2=\sum_{i=9}^{12}i^2 \]

OpenStudy (anonymous):

\[\sum_{k=1}^4(k+8)^2=(1+8)^2+(2+8)^2+(3+8)^2+(4+8)^2\]

OpenStudy (anonymous):

more easily written as \[9^2+10^2+11^2+12^2\] then take out a calculator and add

OpenStudy (anonymous):

so 64,100,121,144

OpenStudy (anonymous):

\[ \sum_{i=9}^{12}i^2 = \sum_{i=0}^{12}i^2 - \sum_{i=0}^{8}i^2 \]

OpenStudy (anonymous):

not 64, 81

OpenStudy (anonymous):

i get 466 http://www.wolframalpha.com/input/?i=9^2%2B10^2%2B11^2%2B12^2

OpenStudy (anonymous):

me too.

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

so the terms are 81,100,121,144 and all that means that equation equals 446